# On robustness of Spectral Rényi divergence

### v3

@@
Abstract
This paper studies a specific category of statistical divergences for spectral densities of time series: the spectral $\alpha$-Rényi divergences, which includes the Itakura–Saito divergence as a subset. The aim of this paper is to highlight both information-theoretic and statistical properties of spectral $\alpha$-Rényi divergences. We reveal the connection between the spectral $\alpha$-Rényi divergence and the $\gamma$-divergence in robust statistics, and a variational representation of spectral $\alpha$-Rényi divergence. Inspired by these results suggesting “robustness” of spectral $\alpha$-Rényi divergence, we show that the minimum spectral Rényi divergence estimate has a stable optimization path with respect to outliers in the frequency domain, unlike the minimum Itakura–Saito divergence estimator, and thus it delivers more stable estimate, reducing the need for intricate pre-processing.
keywords:
Frequency domain analysis; Optimization theory; Robust statistics; Spectral density; Statistical divergence; Time series analysis
and
@@ 1 Introduction
Frequency-domain analysis of time series data has been conducted in many applied fields. Central to this analysis is the (power) spectral density, a crucial element whose estimation has garnered significant attention. For instance, in seismology, earthquake source parameters are delineated through the spectral parameter estimation [5, 33]. In geodesy, temporal correlation of noises in the Global Navigation Satellite System is captured by spectral parameter estimation [17]. In audio signal processing, various methodologies have been developed via spectral density estimation to achieve signal separation [20].
@@
Traditional spectral estimation often hinges on the Whittle likelihood maximization [31], which is equivalent to minimizing the Itakura–Saito divergence [13] between a periodogram and a selected class of spectral densities. Broadening the perspective, one can approach such estimations through the lens of spectral divergences, statistical metrics for dissimilarities between two spectral densities. Estimation based on general spectral divergences is comprehensively explored by [28].
@@
In this paper, we take another look at a particular class of spectral divergence: the spectral $\alpha$-Rényi divergences [29, 34, 32, 8, 9]. Notably, the spectral Rényi divergence encompasses the Itakura–Saito divergence as a special instance. In the probabilistic and information-theoretic literature, this class has been mentioned in the extended discussion of “Pinsker’s information theoretic justification of the Itakura-Saito distortion measure” [24]. In the statistical literature, the robustness of the spectral Rényi divergence against spectral peaks is pointed out [34], which has been utilized in clustering analysis of time series [32, 11].
@@
The primary aim of this paper is to elucidate new properties of this class that shed light on the robustness in the spectral parameter estimation. Specifically, we show that (i) the $\gamma$-divergence [7] between probability densities of the processes leads to the spectral Rényi divergence, and (ii) the spectral Rényi divergence has a variational representation. Inspired by these new results, we explore (iii) an additional robustness property of the minimum spectral Rényi divergence estimate. The minimum spectral Rényi divergence estimate has been shown to be robust against time series outliers in the frequency domain [34, 32, 11]. Expanding on these findings, we show that the optimization path in obtaining the minimum spectral Rényi divergence estimate is stable in the presence of time series outliers, a quality that the minimum Itakura–Saito divergence estimator does not have.
@@
Let us explain the implications of our results. The first result bridges spectral and probabilistic divergences, providing an information-theoretic justification of the spectral Rényi divergence. The second result gives a decomposition of spectral Rényi divergence in terms of the Itakura–Saito divergence. The first two results contribute to understanding the robustness of the spectral Rényi divergence. The third result has both theoretical and practical implications. Theoretically, it gives a novel connection between robust spectral analysis (c.f., [19]) and optimization theory. Developing robust statistics from an algorithmic or computational perspective has been a recent topic of discussion (c.f., [6]). Our results provide an additional perspective on robust statistics from a computational viewpoint, focusing on the stability of optimization paths. Practically, outliers in the frequency domain often emerge due to insufficient detrending [10, 12, 22]. Yet, when using the Itakura–Saito divergence, practitioners should exercise greater caution regarding detrending and other pre-processing steps. In comparison, spectral Rényi divergences offer more stable estimation results without necessitating specific meticulous pre-processing.
@@
The structure of this paper is as follows. Section 2 delivers new formulae related to the spectral Rényi divergence (Theorem 1 and Theorem 2). Section 3 discusses the stability of optimization paths of the spectral Rényi divergence minimization (Theorem 3). Section 4 presents thorough numerical studies. Proofs of the results are presented in Appendices.
@@ 2 Spectral Rényi divergence
This section introduces a class of spectral Rényi divergences and its properties. Let $\mathcal{S}:=\{S(\cdot):\int_{-\pi}^{\pi}S(\omega)\,\mathrm{d}\omega<\infty,\,S(\omega)>0\,\text{for all $\omega\in(0,\pi)$},S(\omega)=S(-\omega)\}$. Fix $S\in\mathcal{S}$ and $\widetilde{S}\in\mathcal{S}$ arbitrarily. Let $p_{n}(\mathbf{x}_{n})$ and $\widetilde{p}_{n}(\mathbf{x}_{n})$, $\mathbf{x}_{n}\in\mathbb{R}^{n}$, be the $n$-th order probability densities of two zero-mean stationary possibly non-Gaussian processes with spectral densities $S$ and $\widetilde{S}$, respectively. We assume the sampling rate is equal to 1.
@@
For $\alpha=1$, the spectral $\alpha$-Rényi divergence is defined by using the left-hand limit:
$D_{1}[\,S\,:\,\widetilde{S}\,]$ $:=\lim_{\alpha\to 1^{-}}D_{\alpha}[\,S\,:\,\widetilde{S}\,]$
$=\frac{1}{2\pi}\int_{-\pi}^{\pi}\left\{\left(\frac{\widetilde{S}(\omega)}{S(\omega)}\right)^{-1}-1+\log\left(\frac{\widetilde{S}(\omega)}{S(\omega)}\right)\right\}\,\mathrm{d}\omega,$ (2)
which is equal to the Itakura–Saito divergence [25, 13].
The name of this class comes from the fact that the class is induced by the limit of the probabilistic Rényi divergence between stationary Gaussian processes, which was pointed out by [29]. That is, if two densities $p_{n}$ and $\widetilde{p}_{n}$ are Gaussian, the following holds for $\alpha\in(0,1]$:
@@
The spectral $\alpha$-Rényi divergence with $\alpha\in(0,1]$ is actually a (statistical) divergence: that is, $D_{\alpha}[S,\widetilde{S}]\geq 0$ and $D_{\alpha}[S,\widetilde{S}]=0$ if and only if $S=\widetilde{S}$. This follows since the one-step ahead prediction error variance (the innovation variance) derived from the Kolmogorov–Szëgo formula (Theorem 5.8.1. of [3])
@@
For $\alpha\in(0,1]$, we shall prepare a discrete version of the spectral Rényi divergence as follows:
@@
where $\Omega_{n}:=\{2\pi(t/n):t=-\lceil n/2\rceil+1,\ldots,-1,0,1,\ldots,\lfloor n/2\rfloor\}$. By the convergence of the Riemann sums, $D^{(n)}_{\alpha}$ converges to $D_{\alpha}$ as $n$ goes to infinity.
@@ 2.1 Properties of spectral Rényi divergences
Here we present two new formulae that have not been investigated in the literature: (i) the asymptotic equivalence between the $\gamma$-divergence [7] and the spectral Rényi divergence; and (ii) the variational representation of spectral Rényi divergences.
We first state the asymptotic equivalence between the $\gamma$-divergence and the spectral Rényi divergence. Let $G_{\gamma}[\,p\,:\,q\,]$ be the $\gamma$-divergence between two probability densities $p$ and $q$:
@@
Theorem 1 (The $\gamma$-divergence leads to spectral Rényi divergences).
Let $\alpha\in(0,1)$ and $S,\widetilde{S}\in\mathcal{S}$. If two densities $p_{n}$ and $\widetilde{p}_{n}$ are the $n$-th order probability densities of stationary Gaussian processes with spectral densities $S$ and $\widetilde{S}$, we have
@@
The proof is given in Appendix A. It employs the direct evaluation of the $\gamma$-divergence and Szegö’s limit theorem[27, 15].
@@
$D_{\alpha}\,[\,S\,:\,\widetilde{S}\,]$ $=\frac{1}{1-\alpha}\min_{S^{\prime}\in\mathcal{S}}\left\{\alpha\,D_{1}\,[\,\widetilde{S}\,:\,S^{\prime}\,]+(1-\alpha)\,D_{1}\,[\,S\,:\,S^{\prime}\,]\right\},$ (3)
@@
$\alpha D_{1}[\,\widetilde{S}\,:\,S^{\prime}\,]+(1-\alpha)D_{1}[\,S\,:\,S^{\prime}\,]$
@@
$=D_{1}[\,\alpha\widetilde{S}+(1-\alpha)S\,:\,S^{\prime}\,]$
@@
$D_{\alpha}\,[\,S\,:\,\widetilde{S}\,]$ $=\frac{1}{1-\alpha}\min_{S^{\prime}\in\mathcal{S}}\left\{\alpha\,D_{1}\,[\,S^{\prime}\,:\,S\,]+(1-\alpha)\,D_{1}\,[\,S^{\prime}\,:\,\widetilde{S}\,]\right\},$ (4)
@@
$\alpha D_{1}[\,S^{\prime}\,:\,S\,]+(1-\alpha)D_{1}[\,S^{\prime}\,:\,\widetilde{S}\,]$
@@
$=D_{1}[\,S^{\prime}\,:\,\overline{S}\,]+(1-\alpha)D_{\alpha}\,[\,S\,:\,\widetilde{S}\,].$
@@
Remark that this variational representation is a spectral version of the characterization of Rényi divergence in relation to the composite hypothesis testing [26, 30].
2.2 Robustness of the spectral Renyi divergence
A tantalizing feature of the spectral $\alpha$-Rényi divergence for $\alpha\in(0,1)$ is that it is robust with respect to outliers in the frequency domain, which has been discussed in [34, 11]. Theorem 1 explains about this robustness as the $\gamma$-divergence induces the robust parameter estimation scheme [7]. Theorem 2 also provides another explanation about the robustness of the spectral Rényi divergence as discussed in this subsection. Let $\mathcal{S}_{\Theta}:=\{S_{\theta}\in\mathcal{S}:\theta\in\Theta\}$ be a parametric spectral model with a parameter space $\Theta\subset\mathbb{R}^{d}$. Consider the minimum spectral $\alpha$-Rényi divergence estimator $\widehat{\theta}_{\alpha}$ given as
@@ 2.2 Robustness of the spectral Renyi divergence
where $\widetilde{I}_{n}$ is a nonparametric pilot estimate of the spectral density such as the periodgram
@@
We begin with introducing outliers and contamination in the frequency domain; see Section 8 of [19] for details. Expressing contamination caused by outliers in the frequency domain, we consider a contaminated pilot estimate $\widetilde{I}_{n}^{z,\omega^{*}}$: for $\omega^{*}\in\Omega_{n}$ and $z>0$,
@@
This contamination is motivated by the case when periodic components are not suitably subtracted and are contaminated in observed time series [10, 12, 22], say, $x_{t}^{o}=x_{t}+\sqrt{8\pi z/n}\cos(At)$, $t=1,\ldots,n$ with original zero-mean time series $\{x_{t}\}_{t=1,2,\cdots,n}$. In this case, the periodogram of $x_{t}^{o}$ has approximately the form of
@@
$D_{1}^{(n)}[\widetilde{I}^{z,\omega^{*}}_{n}\,:\,S_{\theta}]-D_{1}^{(n)}[\widetilde{I}_{n}\,:\,S_{\theta}]=\frac{z}{n}\frac{1}{S_{\theta}(\omega^{*})}-\frac{\log z}{n}+o_{P}(1)$ (5)
@@
$D^{(n)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}\,:\,\alpha S_{\theta}+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}]-D^{(n)}_{1}[\widetilde{I}_{n}\,:\,\alpha S_{\theta}+(1-\alpha)\widetilde{I}_{n}]$

### v4

@@
Abstract: This paper studies a specific class of statistical divergences for spectral densities of time series: the spectral $\alpha$-Rényi divergences, which include the Itakura–Saito divergence as a limiting case. The aim of this paper is to highlight both information-theoretic and statistical properties of spectral $\alpha$-Rényi divergences. We reveal the connection between the spectral $\alpha$-Rényi divergence and the $\gamma$-divergence in robust statistics, and a variational representation of the spectral $\alpha$-Rényi divergence. Inspired by these results suggesting “robustness” of spectral $\alpha$-Rényi divergence, we show that the minimum spectral Rényi divergence estimate has a stable optimization path with respect to outliers in the frequency domain, unlike the minimum Itakura–Saito divergence estimator, and thus it delivers more stable estimates, reducing the need for intricate pre-processing.
Keywords: Frequency domain analysis; Optimization theory; Robust statistics; Spectral density; Statistical divergence; Time series analysis.
@@ 1 Introduction
Frequency-domain analysis of time series data has been conducted in many applied fields. Central to this analysis is the (power) spectral density, a crucial element whose estimation has garnered significant attention. For instance, in seismology, earthquake source parameters are delineated through the spectral parameter estimation Calderoni and Abercrombie (2023); Yoshimitsu et al. (2023). In geodesy, temporal correlation of noises in the Global Navigation Satellite System is captured by spectral parameter estimation Langbein (2004). In audio signal processing, various methodologies have been developed via spectral density estimation to achieve signal separation Martin (2005).
@@
Traditional spectral estimation often hinges on the Whittle likelihood maximization Whittle (1953), which is equivalent to minimizing the Itakura–Saito divergence Itakura and Saito (1968) between a periodogram and a selected class of spectral densities. Broadening the perspective, one can approach such estimations through the lens of spectral divergences, statistical metrics for dissimilarities between two spectral densities. Estimation based on general spectral divergences is comprehensively explored by Taniguchi (1987).
@@
In this paper, we take another look at a particular class of spectral divergences: the spectral $\alpha$-Rényi divergences Vajda (1989); Zhang and Taniguchi (1995); Kakizawa et al. (1998); Gil et al. (2013); Grivel et al. (2021). Notably, a limit of the spectral Rényi divergence induces the Itakura–Saito divergence. In the probabilistic and information-theoretic literature, this class has been mentioned in the extended discussion of “Pinsker’s information-theoretic justification of the Itakura-Saito distortion measure” Parzen (1993). In the statistical literature, the robustness of the spectral Rényi divergence against spectral peaks is pointed out Zhang and Taniguchi (1995), which has been utilized in clustering analysis of time series Kakizawa et al. (1998); Hirukawa (2006).
@@
The primary aim of this paper is to elucidate new properties of this class that shed light on the robustness in the spectral parameter estimation. Specifically, we show that (i) the $\gamma$-divergence Fujisawa and Eguchi (2008) between probability densities of the processes leads to the spectral Rényi divergence, and (ii) the spectral Rényi divergence has a variational representation. Inspired by these new results, we explore (iii) an additional robustness property of the minimum spectral Rényi divergence estimate. The minimum spectral Rényi divergence estimate has been shown to be robust against time series outliers in the frequency domain Zhang and Taniguchi (1995); Kakizawa et al. (1998); Hirukawa (2006). Expanding on these findings, we show that the optimization path in obtaining the minimum spectral Rényi divergence estimate is stable in the presence of time series outliers, a quality that the minimum Itakura–Saito divergence estimator does not have.
@@
Let us explain the implications of our results. The first result bridges spectral and probabilistic divergences, providing an information-theoretic justification of the spectral Rényi divergence. The second result gives a decomposition of spectral Rényi divergence in terms of the Itakura–Saito divergence. The first two results contribute to understanding the robustness of the spectral Rényi divergence. The third result has both theoretical and practical implications. Theoretically, it gives a novel connection between robust spectral analysis (c.f., Maronna et al. (2019)) and optimization theory. Developing robust statistics from an algorithmic or computational perspective has been a recent topic of discussion (c.f., Diakonikolas and Kane (2023)). Our results provide an additional perspective on robust statistics from a computational viewpoint, focusing on the stability of optimization paths. Practically, outliers in the frequency domain often emerge due to insufficient detrending Heyde and Dai (1996); Iacone (2010); McCloskey and Perron (2013). Yet, when using the Itakura–Saito divergence, practitioners should exercise greater caution regarding detrending and other pre-processing steps. In comparison, spectral Rényi divergences offer more stable estimation results without necessitating elaborate pre-processing. This aspect is further supported by an applied study Kano et al. (2025).
@@
The structure of this paper is as follows. Section 2 delivers new formulae related to the spectral Rényi divergence (Theorems 1 and 2). Section 3 discusses the stability of optimization paths of the spectral Rényi divergence minimization (Theorem 3). Section 4 presents thorough numerical studies. Proofs of the results are presented in Appendices.
@@ 2 Spectral Rényi divergence
This section introduces a class of spectral Rényi divergences and their properties. Let $\mathcal{S}:=\{S:[-\pi,\pi]\to[0,\infty]\mid\int_{-\pi}^{\pi}S(\omega)\,\mathrm{d}\omega<\infty,\,S(\omega)>0\,\mbox{and}\,S(\omega)=S(-\omega)\,\text{for all $\omega\in(0,\pi)$}\}$. Fix $S\in\mathcal{S}$ and $\widetilde{S}\in\mathcal{S}$ arbitrarily. Let $p_{n}$ and $\widetilde{p}_{n}$ denote the joint densities of $n$ consecutive observations from two zero-mean stationary processes, possibly non-Gaussian, with spectral densities $S$ and $\widetilde{S}$, respectively. Without loss of generality, we normalize the time scale so that observations are taken at unit time intervals.
@@
For $\alpha=1$, we define $D_{1}[\,S\,:\,\widetilde{S}\,$] by continuity as the limit $\alpha\to 1^{-}$. In this limit, the spectral $\alpha$-Rényi divergence converges to the Itakura–Saito divergence Pinsker (1964); Itakura and Saito (1968):
$\lim_{\alpha\to 1^{-}}D_{\alpha}[\,S\,:\,\widetilde{S}\,$] $=D_{\mathrm{IS}}[\,S\,:\,\widetilde{S}\,$]
$:=\frac{1}{2\pi}\int_{-\pi}^{\pi}\left\{\left(\frac{\widetilde{S}(\omega)}{S(\omega)}\right)^{-1}-1+\log\left(\frac{\widetilde{S}(\omega)}{S(\omega)}\right)\right\}\,\mathrm{d}\omega.$ (2)
The name of this class stems from the fact that the class is induced by the limit of the probabilistic Rényi divergence between stationary Gaussian processes, as pointed out by Vajda (1989). More precisely, if $p_{n}$ and $\widetilde{p}_{n}$ are the joint densities of $n$ consecutive observations of two zero-mean stationary Gaussian processes with spectral densities $S$ and $\widetilde{S}$, respectively, then for $\alpha\in(0,1]$, we have
@@
For $\alpha=1$, the probabilistic Rényi divergence is defined by continuity as the limit $\alpha\to 1^{-}$ and coincides with the Kullback–Leibler divergence:
$D_{1}[\,p_{n}\,:\,\widetilde{p}_{n}\,]=\int_{\mathbb{R}^{n}}p_{n}(\mathbf{x}_{n})\log\frac{p_{n}(\mathbf{x}_{n})}{\widetilde{p}_{n}(\mathbf{x}_{n})}\,\mathrm{d}\mathbf{x}_{n}.$
The spectral $\alpha$-Rényi divergence is actually a (statistical) divergence for any $\alpha\in(0,1)$: that is, $D_{\alpha}[\,S\,:\,\widetilde{S}\,]\geq 0$ and $D_{\alpha}[\,S\,:\,\widetilde{S}\,]=0$ if and only if $S=\widetilde{S}$. This follows since the one-step ahead prediction error variance (the innovation variance) derived from the Kolmogorov–Szegö formula (Theorem 5.8.1. of Brockwell and Davis (1991))
@@
For $\alpha\in(0,1)$, we shall prepare a discrete version of the spectral Rényi divergence as follows:
@@
where $\Omega_{n}:=\{2\pi(t/n):t=-\lceil n/2\rceil+1,\ldots,-1,0,1,\ldots,\lfloor n/2\rfloor\}$. By the convergence of the Riemann sums, $D^{(n)}_{\alpha}$ converges to $D_{\alpha}$ as $n$ goes to infinity. The discrete version $D^{(n)}_{\mathrm{IS}}$ of the Itakura–Saito divergence is defined in the same way.
@@ 2.1 Properties of spectral Rényi divergences
Here we present two new formulae that have not been investigated in the literature: (i) the asymptotic equivalence between the time-domain $\gamma$-divergence Fujisawa and Eguchi (2008) and the spectral Rényi divergence; and (ii) the variational representation of spectral Rényi divergences.
We first state the asymptotic equivalence between the time-domain $\gamma$-divergence and the spectral Rényi divergence. Let $G_{\gamma}[\,p\,:\,q\,]$ be the $\gamma$-divergence between two probability densities $p$ and $q$:
@@
Theorem 1 (The time-domain $\gamma$-divergence leads to the spectral Rényi divergence).
Let $\alpha\in(0,1)$ and $S,\widetilde{S}\in\mathcal{S}$. Let $p_{n}$ and $\widetilde{p}_{n}$ denote the joint densities of $n$ consecutive observations from zero-mean stationary Gaussian processes with spectral densities $S$ and $\widetilde{S}$, respectively. Then, we have
@@
The proof is given in Appendix A. It employs the direct evaluation of the $\gamma$-divergence and Szegö’s limit theorem Szegö (1920); Kolmogorov (1941).
@@
$D_{\alpha}\,[\,S\,:\,\widetilde{S}\,$] $=\frac{1}{1-\alpha}\min_{S^{\prime}\in\mathcal{S}}\left\{\alpha\,D_{\mathrm{IS}}\,[\,\widetilde{S}\,:\,S^{\prime}\,]+(1-\alpha)\,D_{\mathrm{IS}}\,[\,S\,:\,S^{\prime}\,]\right\},$ (3)
@@
$\alpha D_{\mathrm{IS}}[\,\widetilde{S}\,:\,S^{\prime}\,]+(1-\alpha)D_{\mathrm{IS}}[\,S\,:\,S^{\prime}\,]$
@@
$=D_{\mathrm{IS}}[\,\alpha\widetilde{S}+(1-\alpha)S\,:\,S^{\prime}\,]$
@@
[…]
@@
$\alpha D_{\mathrm{IS}}[\,S^{\prime}\,:\,S\,]+(1-\alpha)D_{\mathrm{IS}}[\,S^{\prime}\,:\,\widetilde{S}\,$]
@@
$=D_{\mathrm{IS}}[\,S^{\prime}\,:\,\overline{S}\,]+(1-\alpha)D_{\alpha}\,[\,S\,:\,\widetilde{S}\,$].
@@
Note that this variational representation is a spectral version of the characterization of Rényi divergence in relation to the composite hypothesis testing Shayevitz (2010); van Erven and Harremoës (2014).
2.2 Robustness of the spectral Rényi divergence
A tantalizing feature of the spectral $\alpha$-Rényi divergence is that it is robust with respect to outliers in the frequency domain, which has been discussed in Zhang and Taniguchi (1995); Hirukawa (2006). Theorem 1 explains this robustness as the time-domain $\gamma$-divergence induces the robust parameter estimation scheme Fujisawa and Eguchi (2008). Theorem 2 also provides another explanation about the robustness of the spectral Rényi divergence as discussed in this subsection. Let $\mathcal{S}_{\Theta}:=\{S_{\theta}\in\mathcal{S}:\theta\in\Theta\}$ be a parametric spectral model with a parameter space $\Theta\subset\mathbb{R}^{d}$, where $\Theta$ is compact with nonempty interior.
Consider the minimum spectral $\alpha$-Rényi divergence estimator $\widehat{\theta}_{\alpha}$ given as
@@ 2.2 Robustness of the spectral Rényi divergence
where $\widetilde{I}_{n}$ is a nonparametric pilot estimate of the spectral density such as the periodogram
@@
We begin with introducing outliers and contamination in the frequency domain; see Section 8 of Maronna et al. (2019) for details. Expressing contamination caused by outliers in the frequency domain, we consider a contaminated pilot estimate $\widetilde{I}_{n}^{z,\omega^{*}}$: for $\omega^{*}\in\Omega_{n}$ and $z>0$,
@@
This contamination is motivated by the case when periodic components are not suitably subtracted and are contaminated in observed time series Heyde and Dai (1996); Iacone (2010); McCloskey and Perron (2013), say, $x_{t}^{o}=x_{t}+\sqrt{8\pi z/n}\cos(At)$, $t=1,\ldots,n$ with original zero-mean time series $\{x_{t}\}_{t=1,2,\cdots,n}$. In this case, the periodogram of $x_{t}^{o}$ has approximately the form of
@@
$D_{\mathrm{IS}}^{(n)}[\widetilde{I}^{z,\omega^{*}}_{n}\,:\,S_{\theta}]-D_{\mathrm{IS}}^{(n)}[\widetilde{I}_{n}\,:\,S_{\theta}]=\frac{z}{n}\frac{1}{S_{\theta}(\omega^{*})}-\frac{\log z}{n}+o_{P}(1)$ (5)
@@
$D_{\mathrm{IS}}^{(n)}[\widetilde{I}^{z,\omega^{*}}_{n}\,:\,\alpha S_{\theta}+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}]-D_{\mathrm{IS}}^{(n)}[\widetilde{I}_{n}\,:\,\alpha S_{\theta}+(1-\alpha)\widetilde{I}_{n}]$